Use substitution to solve the system of equations
Y=4x-1
Y=2x-5

Answer by josgarithmetic(39625)

Two formulas for Y in your system; equate then and you have one equation in just the variable, x. Each equation is already solved for Y in terms of just x.
4x-1=Y, so substitute for Y from the second equation.
4x-1=2x-5;
4x-1-2x=-5
2x-1=-5
2x=-5+1
2x=-4
x=-2

Now use either equation to get the value for Y.